Solve the equation x + 6 = √(17x + 32)
Answer by mathjames(8)

x + 6 = √(17x + 32)
{x + 6}square it = {√(17x + 32)}square it also
x {raise to 2}+12x+36=17x+32
x {raise to 2}=17x-12x+32-36
=5x+4
4x=4
x=1